OpenStudy (christos):

I have an equation of a line. My known parts are x1 x2 y1 y2. How do I find "m" and "c"?

OpenStudy (angela210793):

the equation of a line is written as y=mx+c where m=slope and c= y-intercept m=(y2-y1)/(x2-x1) and to find c all u have to do is take x=0

OpenStudy (christos):

@angela210793 I came across this on the internet: 1)m = (y2 - y1) / (x2 - x1) 2)C = y1 - (x1 * m) How can I get to each of the above?

OpenStudy (perl):

hello

OpenStudy (perl):